You are here: Foswiki>Software Web>SWTestingStatus>SWTestReports>BP8 (14 Feb 2011, NormCushing)Edit Attach

BP8 Test Report

This is the high level status report for the TCS BP8 testing.

Because we were not able to observe during the day to perform some basic regression testing version BP8 will not be used for science observing.

So science observing should continue with...

- AO5 for monocular and/or master/slave observing with LBC. - BP7 for Gregorian instrument observing.

On Wednesday night...

We performed a series of closed dome offsets. Each series consisted of 20 dice 5 patterns consisting of absolute and relative RA/DEC offsets. These were performed to help identify a PCS tuning value. The GCS was in full simulation mode so the AGw probe was not exercised during any of these tests.

We also tested binocular Gregorian Presets, Offsets, and guiding. We learned the magnitude of the tip/tilts needed for guiding is reasonable. We also learned the locking/unlocking logic is not correct. Overall this was a big step forward.

On Thursday night...

We used the binocular version of the LBC software to perform monocular guiding on each side and to perform binocular operations.

Before testing, Andrea warned us he had found a bug in the LBC binocular code that affected fits files. We decided to move forward with testing.

Michele said, "We were able to perform Monocular guiding with LBC on both the Blue and Red sides (finally) with BP8. Indeed, the LBC FITS images are corrupted in that they are non-conformant FITS files (i.e., there is pixel data, but the header is corrupted and missing required elements). However, Dave was able to display the image portion and make measurements.

The Blue side had ~1" FWHM and the Red had 0.83" FWHM. Dave and John indicated this was in line with what is currently seen with LBC. In fact, the Red was rather good with little coma. The Blue did have significantly more coma (elongated but not trailed images). Dave felt the guiding was working properly. I will plot up the guide centroids with MatLab when I have a chance.

We tried Binocular Guiding, but this is really a no-go as the locking/unlocking has a logic hole. We "guided" and even seemingly well with the Red, but the Blue was bad (larger and larger tips/tilts). In any case the tip/tilts are now the right order of magnitude and not crazy values."

Chris Biddick said, "For the BP8 LBC testing, we changed the lbccontrol pointer to Binocular, and we also changed the lbccontrol TCS pointer to BP8. Then we rebuilt the LBC control program to insure everything was proper.

It all worked correctly, with the exception of the FITS headers."

Presets were issued. Chris does not recall issuing Offsets.

Chris also said, "When we were finished, we put the lbccontrol pointer back to MasterSlave, put the lbccontrol TCS pointer back to AO2, switched our build to AO5 and tested. I had to edit the lbt.conf file on lbccontrol to change the IIF IP address. After that, everything was fine, including the FITS headers (as expected)."

The IP address change was necessary because as part of the network changes this week we changed the IP subnet for the "tcs" and "obs" computers.

-- NormCushing - 14 Feb 2011
Topic revision: r1 - 14 Feb 2011, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback